OpenStudy (anonymous):

What is the similarity ratio of two cubes with volumes of 64 cubic centimeters and 125 cubic centimeters? 2:3 4:5 1:5 16:25

OpenStudy (anonymous):

\[\large \frac{a_1^3}{a_2^3} = \frac{64}{125} \implies \frac{a_1}{a_2} = \frac{4}{5}\]